# R u s s i a n   A c a d e m y   o f   S c i e n c e s  S i b e r i a n    B r a n c h

### v37

@@ 15.16. Do there exist groups whose rate of growth is $e^{\sqrt{n}}$
a) Such an example is found; moreover, for any countable ordinal $\alpha$ there are groups in $\textbf{H}_{\alpha+1}{\mathfrak{F}}$ that are not in ${\bf H}_{\alpha}{\mathfrak{F}}$ (T. Januszkiewicz, P. H. Kropholler, I. J. Leary, Bull. London Math. Soc., 42, No. 5 (2010), 896–904).
@@
Yes, it exists (N. Dunfield, Appendix B  in S. Kionke, J. Raimbault, Doc. Math., 21 (2016), 873–915).
@@
The set $S$ is finite and consists of 2 and 3. This follows from the complete list of finite simple groups $G$ with a module $V$ in characteristic $p$ satisfying those properties, which was determined in a number of papers culminating in (M. A. Grechkoseeva, S. V. Skresanov, Sibirsk. Elektron. Matem. Izv., 17 (2020), 585–589). Then the group $G$ must be either $3D_{4}(2)$ with $p=2$ (I. B. Gorshkov, Commun. Algebra, 47, no. 12 (2019), 5192–5206), or $U_{5}(2)$ with $p=3$ (A. A. Buturlakin, Algebra Logic, 47, no. 2 (2008), 91–99, and A. V. Zavarnitsine, Siberian Math. J., 49, no. 2 (2008), 246–256), or $U_{3}(q)$, where $q$ is a special Mersenne prime, with $p=2$ (A. V. Zavarnitsine, Algebra Logic, 45, no. 2 (2006), 106–116).
@@
Is it true that $r_{G}(n)$ may be bounded as a function of $\lambda(n)$, the total number (counting multiplicities) of prime factors of $n$? Peter M. Neumann
15.65. A square matrix is said to be separable if its minimal polynomial has no repeated roots, and cyclic if its minimal and characteristic polynomials are equal. For a matrix group $G$ over a finite field define $s(G)$ and $c(G)$ to be the proportion of separable and of cyclic elements respectively in $G$. For a classical group ${\rm X}(d,q)$ of dimension $d$ defined over the field with $q$ elements let  $S(\hbox{\rm X};q):=\lim_{d\to\infty}s(\hbox{\rm X}(d,q))$  and  $C(\hbox{\rm X};q):=\lim_{d\to\infty}c(\hbox{\rm X}(d,q)).$ Independently G. E. Wall (Bull. Austral. Math. Soc., 60, no. 2 (1999), 253–284) and J. Fulman (J. Group Theory, 2, no. 3 (1999), 251–289) have evaluated $S(\hbox{\rm GL};q)$ and $C(\hbox{\rm GL};q)$, and have found them to be rational functions of $q$. Are $S(\hbox{\rm X};q)$ and $C(\hbox{\rm X};q)$ rational functions of $q$ also for the unitary, symplectic, and orthogonal groups? Peter M. Neumann
@@
The groups $SL_{n}({\mathbb{Z}})$, $n>13$, satisfy this condition (M. C. Tamburini, P. Zucca, J. Algebra, 195, no. 2 (1997), 650–661). The groups $PSL_{n}({\mathbb{Z}})$ satisfy it if and only if $n>4$ (N. Ya. Nuzhin, Vladikavkaz. Mat. Zh., 10, no. 1 (2008), 68–74 (Russian)). The group $PSp_{4}({\mathbb{Z}})$ does not satisfy it, which follows from the corresponding fact for $PSp_{4}(3)$, see Archive, 7.30. Ya. N. Nuzhin
@@
Yes, it exists, as follows from Theorem 3 in (I. Z. Ruzsa, Proc. Conf. Number theory (Budapest, 1987). Vol. I: Elementary and analytic, Colloq. Math. Soc. János Bolyai, 51, North-Holland, Amsterdam, 1990, 473–504). Another solution is in (S. V. Skresanov, Siberian Math. J., 61, no. 3 (2020), 542–544).
@@
16.5. A group is said to be perfect if it coincides with its derived subgroup. Does there exist a perfect locally finite $p\thinspace$-group
@@
b) all of whose proper subgroups are solvable?
c) all of whose proper subgroups are hypercentral and solvable? A. O. Asar
@@ B. Poizat
16.88. (G. M. Bergman). The width of a group $G$ with respect to a generating set $X$ means the supremum over all $g\in G$ of the least length of a group word in elements of $X$ expressing $g$. A group $G$ has finite width if the width of $G$ with respect to any generating set is finite. Does there exist a countably infinite group of finite width? All known infinite groups of finite width (infinite permutation groups, infinite-dimensional general linear groups, and some other groups) are uncountable. V. Tolstykh
@@ V. Tolstykh
No, it must not (I. Snopce, S. Tanushevski, P. Zalesskii, Int. Math. Res. Notices, 2022, no. 11 (2022), 8280–8294).
@@ 17.28. Is there a soluble right-orderable group with insoluble word problem?
17.32. Is the following analogue of the Cayley–Hamilton theorem true for the free group $F_{n}$ of rank $n$: If $w\in F_{n}$ and $\varphi\in{\rm Aut\,}F_{n}$ are such that $\langle w,\varphi(w),\dots,\varphi^{n}(w)\rangle=F_{n}$, then $\langle w,\varphi(w),\dots,\varphi^{n-1}(w)\rangle=F_{n}$? O. V. Bogopol’skiĭ
@@
Editors’ comments: Reduction of part (b) to almost simple group $G$ is obtained in (E. P. Vdovin, J. Algebra Appl., 11, no. 1 (2012), 1250015 (14 pages)). An affirmative answer to (b) has been obtained for almost simple groups with socle isomorphic to an alternating group (A. A. Baikalov, Algebra Logic, 56 (2017), 87–97), to a sporadic simple group (T. C. Burness, Israel J. Math., 254 (2023), 313–340), to $PSL(n,q)$ (A. A. Baykalov, J. Group Theory, 28 (2025), 1003–1077),  to $PSU(n,q)$ or $PSp(n,q)^{\prime}$ (A. A. Baykalov, Int. J. Algebra Comput., 35, no. 06 (2025), 823–908). E. P. Vdovin
@@
a) every subgroup is a $D_{\pi}$-group (H. Wielandt);
b) every subgroup possessing a Hall $\pi$-subgroup is a $D_{\pi}$-group.
All finite simple $D_{\pi}$-groups are known (D. O. Revin, Algebra and Logic, 47, no. 3 (2008), 210–227). Comment of 2013: alternating and sporadic groups for both parts are found in (N. Ch. Manzaeva, Siberian Electr. Math. Rep., 9 (2012), 294–305 (Rus-sian)). E. P. Vdovin, D. O. Revin
@@
17.61. The group ${\rm CT}_{\mathcal{P}}(\mathbb{Z})$ is finitely generated if and only if $\mathcal{P}$ is finite. If $\mathcal{P}=\nobreak\varnothing$, then it is isomorphic to the finitely presented (first) Higman–Thompson group (J. P. McDermott, see Remark 1.4 in (S. Kohl, J. Group Theory, 20, no. 5 (2017), 1025–1030). Is it always finitely presented if $\mathcal{P}$ is finite? S. Kohl
@@
17.84. An associative algebra $A$ is said to be Calabi–Yau of dimension $d$ (for short, CYdd) if there is a natural isomorphism of $A$-bimodules ${\rm Ext}_{A\text{-bimod}}^{d}(A,\,A\otimes A)\cong\nobreak A$ and ${\rm Ext}_{A\text{-bimod}}^{n}(A,\,A\otimes A)=0$ for $n\neq d$. By Kontsevich’s theorem, the complex group algebra ${\mathbb{C}}G$ of the fundamental group $G$ of a 3-dimensional aspherical manifold is CY33. Is every group with CY33 complex group algebra residually finite? R. Mikhailov
@@
b) For which finite simple groups $S$ are the groups $S\times S$ acting on $S$ by $(g,h):x\mapsto g^{-1}xh$ non-synchronizing? Peter M. Neumann
@@
Peter M. Neumann
17.97. Is every variety of groups of exponent $4$ finitely based? Peter M. Neumann
@@
b) Is it true that a variety is small if and only if all its finitely generated groups have the Hopf property? Peter M. Neumann
@@
17.103. Does there exist a continuum of sets $\pi$ of primes for which every finite group possessing a Hall $\pi$-subgroup is a $D_{\pi}$-group? D. O. Revin
@@ 18.6. Is every finitely generated one-relator group residually amenable?
This is the strongest form of the conjecture. We would be also interested in a proof that this holds for $n$ sufficiently large. The result for $r\leqslant 2$, which includes the first three cases above, is proved for $n$ odd (D. Bubboloni, C. E. Praeger, J. Combin. Theory (A), 118 (2011), 2000–2024). When $r\geqslant 3$ we know that $cn\leqslant\gamma(S_{n})\leqslant\frac{2}{3}n$ for some positive constant $c$ (D. Bubboloni, C. E. Praeger, P. Spiga, J. Algebra, 390 (2013) 199–215). We showed that the conjectured value for $\gamma(S_{n})$ is an upper bound, by constructing a normal covering for $S_{n}$ with this number of conjugacy classes of maximal subgroups, and gave further evidence for the truth of the conjecture in other cases (D. Bubboloni, C. E. Praeger, P. Spiga, Int. J. Group Theory, 3, no. 2 (2014), 57–75).
A. Maróti showed that the conjecture is incorrect for odd $n$; his counterexample is presented in § 2 of (D. Bubboloni, C. E. Praeger, P. Spiga, Monatsh. Math., 191 (2020), 229–247). D. Bubboloni, C. E. Praeger, P. Spiga
@@
No, it is not true, for example, for $G$ being an extension of $L_{2}(27)$ by a field automorphism of order $3$, in which $H$ is an extension of a Borel subgroup of $L_{2}(27)$ by a field automorphism of order $3$ (V. I. Zenkov, Abstracts of Int. Conf. “Mal’tsev Meeting 2020”, Novosibirsk, 2020, p. 149).
@@
No, there exists an uncountable connected Polish group whose every abstract homomorphism into $S$ is trivial (A. Kaïchouh, F. Le Maître, Bull. London Math. Soc., 47 (2015), 996–1009). (S. Corson, Letter of 5 June 2024).
@@ 18.70. Is every finitely generated Coxeter group conjugacy separable?
18.89. Consider the set of balanced presentations $\langle x_{1},\dots,x_{n}\mid r_{1},\dots,r_{n}\rangle$ with fixed $n$ generators $x_{1},\dots,x_{n}$. By definition, the group ${AC}_{n}$ of Andrews–Curtis moves on this set of balanced presentations is generated by the Nielsen transformations together with conjugations of relators. Is ${AC}_{n}$ finitely presented?
Comment of 2023: It has been proved that ${AC}_{2}$ is not finitely presented (V. A. Roman’kov, Preprint, 2023, https://arxiv.org/pdf/2305.11838.pdf). J. Swan, A. Lisitsa
@@
The question of which finite simple groups are (2,3)-generated remains open only for the orthogonal groups (M. A. Pellegrini, M. C. Tamburini Bellani, J. Algebra, 598 (2022), 156–193). M. C. Tamburini
@@
18.102. (E. C. Dade). Let $C$ be a Carter subgroup of a finite solvable group $G$, and let $\ell(C)$ be the number of primes dividing $|C|$ counting multiplicities. It was proved in (E. C. Dade, Illinois J. Math., 13 (1969), 449–514) that there is an exponential function $f$ such that the nilpotent length of $G$ is at most $f(\ell(C))$. Is there a linear (or at least a polynomial) function $f$ with this property? A. Turull
@@
18.104. Construct an example of a pro-$p$ group $G$ and a proper abstract normal subgroup $K$ such that $G/K$ is perfect. J. Wilson
@@
If so, then ${\rm SL}_{3}(R)$ and the kernel of the map to ${\rm SL}_{3}(R/U)$ provide an answer to the previous question. J. Wilson
@@
The motivation is the theorem in (E. Breulliard, B. Green, M. Kassabov, A. Lubotzky, N. Nikolov, T. Tao) saying that there are $\epsilon>0$ and $k$ such that every finite simple group $G$ contains a generating set $S$ with at most $k$ elements such that the Cayley graph of $G$ with respect to $S$ is an $\epsilon$-expander. In (M. Ershov, A. Jaikin-Zapirain, M. Kassabov, Mem. Amer. Math. Soc., 1186, 2017) it is proved that there exists a group with Kazhdan’s property (T) that maps onto every simple group of Lie type of rank $\geqslant 2$. (A group $G$ is said to have Kazhdan’s property (T) if its trivial representation is an isolated point in the unitary dual space of $G$.) A. Jaikin-Zapirain
@@
Road closure conjecture: these necessary conditions are sufficient for the road closure property. P. J. Cameron
@@
The conjecture is proved if $\varphi$ has prime order (E. Jabara, J. Algebra, 320, no. 10 (2008), 3671–3679). If $G$ is infinitely generated, then $G$ does not have to be almost solvable (K. Dekimpe, D. Gonçalves, Bull. London Math. Soc., 46, no. 4 (2014), 737–746). A. L. Fel’shtyn, E. V. Troitsky
@@
19.31. Let $\omega(G,S)$ be the exponential growth rate of a group $G$ with a finite generating set $S$ (see 14.7). Let $MCG(\Sigma_{g})$ be the mapping class group of the orientable surface $\Sigma_{g}$ of genus $g$. Is there a constant $C>1$ such that $\omega(MCG(\Sigma_{g}),S)\geqslant C$ for every $g>0$ and every finite generating set $S$ of $MCG(\Sigma_{g})$?
If the genus $g$ is fixed, then such a constant $C>1$ does exist (J. Mangahas, Geom. Funct. Anal. 19, no. 5 (2010), 1468–1480). K. Fujiwara

### v38

@@ 15.16. Do there exist groups whose rate of growth is $e^{\sqrt{n}}$
Yes it is true; moreover, for any two nilpotent subgroups $H,K$ of a (non-abelian) finite simple group $G$ there is $g\in G$ such that $H\cap K^{g}=1$ (T. C. Burness, H. Y. Huang, Preprint, 2025, https://arxiv.org/abs/2508.03479).
@@
a) Such an example is found; moreover, for any countable ordinal $\alpha$ there are groups in $\textbf{H}_{\alpha+1}{\mathfrak{F}}$ that are not in ${\bf H}_{\alpha}{\mathfrak{F}}$ (T. Januszkiewicz, P. H. Kropholler, I. J. Leary, Bull. London Math. Soc., 42, No. 5 (2010), 896–904). Furthermore, torsion-free examples with similar properies have been constructed (F. Fournier-Facio, B. Sun, Preprint, 2025, https://arxiv.org/abs/2503.01987v3).
@@
Yes, there does (N. Dunfield, Appendix B  in S. Kionke, J. Raimbault, Doc. Math., 21 (2016), 873–915).
@@
The set $S$ is finite and consists of 2 and 3. This follows from the complete list of finite simple groups $G$ with a module $V$ in characteristic $p$ satisfying those properties, which was determined in a number of papers, the last of which is (M. A. Grechkoseeva, S. V. Skresanov, Sibirsk. Elektron. Matem. Izv., 17 (2020), 585–589). Then the group $G$ must be either $3D_{4}(2)$ with $p=2$ (V. D. Mazurov, Algebra Logic, 52, no. 5 (2013), 400–403), or $U_{5}(2)$ with $p=3$ (A. V. Zavarnitsine, Siberian Math. J., 49, no. 2 (2008), 246–256, and M. A. Grechkoseeva, J. Algebra, 339, no. 1 (2011), 304–319), or $U_{3}(q)$, where $q$ is a special Mersenne prime, with $p=2$ (A. V. Zavarnitsine, Algebra Logic, 45, no. 2 (2006), 106–116).
@@
Is it true that $r_{G}(n)$ may be bounded as a function of $\lambda(n)$, the total number (counting multiplicities) of prime factors of $n$? P. M. Neumann
15.65. A square matrix is said to be separable if its minimal polynomial has no repeated roots, and cyclic if its minimal and characteristic polynomials are equal. For a matrix group $G$ over a finite field define $s(G)$ and $c(G)$ to be the proportion of separable and of cyclic elements respectively in $G$. For a classical group ${\rm X}(d,q)$ of dimension $d$ defined over the field with $q$ elements let $S(\hbox{\rm X};q):=\lim_{d\to\infty}s(\hbox{\rm X}(d,q))$ and $C(\hbox{\rm X};q):=\lim_{d\to\infty}c(\hbox{\rm X}(d,q)).$ Independently G. E. Wall (Bull. Austral. Math. Soc., 60, no. 2 (1999), 253–284) and J. Fulman (J. Group Theory, 2, no. 3 (1999), 251–289) have evaluated $S(\hbox{\rm GL};q)$ and $C(\hbox{\rm GL};q)$, and have found them to be rational functions of $q$. Are $S(\hbox{\rm X};q)$ and $C(\hbox{\rm X};q)$ rational functions of $q$ also for the unitary, symplectic, and orthogonal groups? P. M. Neumann
@@
The groups $SL_{n}({\mathbb{Z}})$, $n>13$, satisfy this condition (M. C. Tamburini, P. Zucca, J. Algebra, 195, no. 2 (1997), 650–661). The groups $PSL_{n}({\mathbb{Z}})$ satisfy it if and only if $n>4$ (N. Ya. Nuzhin, Vladikavkaz. Mat. Zh., 10, no. 1 (2008), 68–74 (Russian)). The group $PSp_{4}({\mathbb{Z}})$ does not satisfy it, which follows from the corresponding fact for $PSp_{4}(3)$, see Archive, 7.30. The group $G_{2}$ satisfies it (I. A. Timofeenko, J. Siberian Fed. Univ. Ser. Math. Phys., 8, no. 1 (2015), 104–108. The group $E_{l}$ satisfies it (I. A. Timofeenko, Sibrsk. Èlectron. Mat. Izv., 14 (2017), 807–820 (Russian)).
Ya. N. Nuzhin
@@
Yes, there does, as follows from Theorem 3 in (I. Z. Ruzsa, Proc. Conf. Number theory (Budapest, 1987). Vol. I: Elementary and analytic, Colloq. Math. Soc. János Bolyai, 51, North-Holland, Amsterdam, 1990, 473–504). Another solution is in (S. V. Skresanov, Siberian Math. J., 61, no. 3 (2020), 542–544).
@@
16.5. A group is said to be perfect if it coincides with its derived subgroup. Does there exist a perfect locally finite $p$-group
@@
b) all of whose proper subgroups are solvable? A. O. Asar
@@ B. Poizat
16.88. (G. M. Bergman). The width of a group $G$ with respect to a generating set $X$ means the supremum over all $g\in G$ of the least length of a group word in elements of $X$ expressing $g$. A group $G$ has finite width if the width of $G$ with respect to every generating set is finite. Does there exist a countably infinite group of finite width?
All known infinite groups of finite width (infinite permutation groups, infinite-dimensional general linear groups, and some other groups) are uncountable (G. M. Bergman, Bull. London Math. Soc., 38 (2006), 429–440, and references therein). V. Tolstykh
@@ V. Tolstykh
No, it need not (I. Snopce, S. Tanushevski, P. Zalesskii, Int. Math. Res. Notices, 2022, no. 11 (2022), 8280–8294).
@@ 17.28. Is there a soluble right-orderable group with insoluble word problem?
17.32. Is the following analogue of the Cayley–Hamilton theorem true for the free group $F_{n}$ of rank $n$: If $w\in F_{n}$ and $\varphi\in{\rm Aut\,}F_{n}$ are such that $\langle w,\varphi(w),\dots,\varphi^{n}(w)\rangle=F_{n}$, then $\langle w,\varphi(w),\dots,\varphi^{n-1}(w)\rangle=F_{n}$? O. V. Bogopolski
@@
Editors’ comments: Reduction of part (b) to almost simple group $G$ is obtained in (E. P. Vdovin, J. Algebra Appl., 11, no. 1 (2012), 1250015 (14 pages)). An affirmative answer to (b) has been obtained for almost simple groups with socle isomorphic to an alternating group (A. A. Baikalov, Algebra Logic, 56 (2017), 87–97), to a sporadic simple group (T. C. Burness, Israel J. Math., 254 (2023), 313–340), to $PSL(n,q)$ (A. A. Baykalov, J. Group Theory, 28 (2025), 1003–1077), to $PSU(n,q)$ or $PSp(n,q)^{\prime}$ (A. A. Baykalov, Int. J. Algebra Comput., 35, no. 06 (2025), 823–908). Part (b) also has an affirmative answer if $S$ is a maximal subgroup of $G$ (T. C. Burness, Algebra Number Theory, 15, no. 7 (2021), 1755–1807). E. P. Vdovin
@@
(a) every subgroup is a $D_{\pi}$-group (H. Wielandt);
(b) every subgroup possessing a Hall $\pi$-subgroup is a $D_{\pi}$-group.
All finite simple $D_{\pi}$-groups are known (D. O. Revin, Algebra Logic, 47, no. 3 (2008), 210–227). Comment of 2013: Alternating and sporadic groups for both parts are found in (N. Ch. Manzaeva, Siberian Electron. Math. Rep., 9 (2012), 294–305 (Russian)). Comment of 2025: Simple groups of Lie type of rank 1 satisfying condition (a) are known (D. O. Revin, V. D.  Shepelev, Siberian Math. J., 65, no. 5 (2024), 1187–1194, V. D.  Shepelev, Siberian Math. J., 66, no. 4 (2025), 1049–1062). E. P. Vdovin, D. O. Revin
@@
17.61. The group ${\rm CT}_{\mathcal{P}}(\mathbb{Z})$ is finitely generated if and only if $\mathcal{P}$ is finite. If $\mathcal{P}=\penalty 10000\varnothing$, then it is isomorphic to the finitely presented (first) Higman–Thompson group (J. P. McDermott, see Remark 1.4 in S. Kohl, J. Group Theory, 20, no. 5 (2017), 1025–1030). Is it always finitely presented if $\mathcal{P}$ is finite? S. Kohl
@@
Yes, such a group exists, and there are infinitely many examples (S. V. Skresanov, Preprint, 2025, https://arxiv.org/abs/2509.17587).
@@
17.84. An associative algebra $A$ is said to be Calabi–Yau of dimension $d$ (for short, CYdd) if there is a natural isomorphism of $A$-bimodules ${\rm Ext}_{A\text{-bimod}}^{d}(A,\,A\otimes A)\cong\penalty 10000A$ and ${\rm Ext}_{A\text{-bimod}}^{n}(A,\,A\otimes A)=0$ for $n\neq d$. By Kontsevich’s theorem, the complex group algebra ${\mathbb{C}}G$ of the fundamental group $G$ of a 3-dimensional aspherical manifold is CY33. Is every group with CY33 complex group algebra residually finite? R. Mikhailov
@@
b) For which finite simple groups $S$ are the groups $S\times S$ acting on $S$ by $(g,h):x\mapsto g^{-1}xh$ non-synchronizing? P. M. Neumann
@@
P. M. Neumann
17.97. Is every variety of groups of exponent $4$ finitely based? P. M. Neumann
@@
b) Is it true that a variety is small if and only if all its finitely generated groups have the Hopf property? P. M. Neumann
@@
17.103. Does there exist a continuum of sets $\pi$ of primes for which every finite group possessing a Hall $\pi$-subgroup is a $D_{\pi}$-group?
Comment of 2025: It has been proved that, for any integer $x$, if $\pi$ is the set of all prime numbers $p>x$, then any finite group containing a Hall $\pi$-subgroup is a $D_{\pi}$-group (K. A. Ilenko, N. V. Maslova, Siberian Math. J., 62, no. 1 (2021), 44–51). D. O. Revin
@@ 18.6. Is every finitely generated one-relator group residually amenable?
This is the strongest form of the conjecture. We would be also interested in a proof that this holds for $n$ sufficiently large. The result for $r\leqslant 2$, which includes the first three cases above, is proved for $n$ odd (D. Bubboloni, C. E. Praeger, J. Combin. Theory (A), 118 (2011), 2000–2024). When $r\geqslant 3$ we know that $cn\leqslant\gamma(S_{n})\leqslant\frac{2}{3}n$ for some positive constant $c$ (D. Bubboloni, C. E. Praeger, P. Spiga, J. Algebra, 390 (2013) 199–215). We showed that the conjectured value for $\gamma(S_{n})$ is an upper bound, by constructing a normal covering for $S_{n}$ with this number of conjugacy classes of maximal subgroups, and gave further evidence for the truth of the conjecture in other cases (D. Bubboloni, C. E. Praeger, P. Spiga, Int. J. Group Theory, 3, no. 2 (2014), 57–75). D. Bubboloni, C. E. Praeger, P. Spiga
A. Maróti showed that the conjecture is incorrect for odd $n$; his counterexample is presented in § 2 of (D. Bubboloni, C. E. Praeger, P. Spiga, Monatsh. Math., 191 (2020), 229–247).
@@
No, it is not true, for example, for $G$ being an extension of $L_{2}(27)$ by a field automorphism of order $3$, in which $H$ is an extension of a Borel subgroup of $L_{2}(27)$ by a field automorphism of order $3$ (V. I. Zenkov, Siberian Math. J., 63, no. 4 (2022), 720–722).
@@
No, there exists an uncountable connected Polish group whose every abstract homomorphism into $S$ is trivial (C. Rosendal, S. Solecki, Israel J. Math., 162 (2007), 349–371); see also (A. Kaïchouh, F.  Le Maître, Bull. London Math. Soc., 47 (2015), 996–1009). (S. Corson, Letter of 5 June 2024).
@@ 18.70. Is every finitely generated Coxeter group conjugacy separable?
18.89. Consider the set of balanced presentations $\langle x_{1},\dots,x_{n}\mid r_{1},\dots,r_{n}\rangle$ with fixed $n$ generators $x_{1},\dots,x_{n}$. By definition, the group ${AC}_{n}$ of Andrews–Curtis moves on this set of balanced presentations is generated by the Nielsen transformations together with conjugations of relators. Is ${AC}_{n}$ finitely presented? J. Swan, A. Lisitsa
The group ${AC}_{2}$ is not finitely presented (S. Krstić, J. McCool, J. London Math. Soc. (2), 56, no. 2 (1997), 264–274; also V. A. Roman’kov, Preprint, 2023, https://arxiv.org/pdf/2305.11838.pdf), while the groups ${AC}_{n}$, $n\geqslant 3$, are finitely presented, which follows from (G. Kiralis, S. Krstič, J. McCool, Proc. London Math. Soc. (3), 73, no. 3 (1996), 481–720). (M. Ershov, Letter of 16 October 2025, https://kourovkanotebookorg.wordpress.com/wp-content/uploads/2025/11/kourovka1889.pdf).
@@
The question of which finite simple groups are (2,3)-generated remains open only for the orthogonal groups of even dimension $2m>8$ (M. A. Pellegrini, M. C. Tamburini Bellani, J. Austral. Math. Soc., 117 (2024), 130–148). M. C. Tamburini
@@
18.102. (E. C. Dade). Let $C$ be a Carter subgroup of a finite solvable group $G$, and let $\ell(C)$ be the number of primes dividing $|C|$ counting multiplicities. It was proved in (E. C. Dade, Illinois J. Math., 13 (1969), 449–514) that there is an exponential function $f$ such that the nilpotent length of $G$ is at most $f(\ell(C))$. Is there a linear (or at least a polynomial) function $f$ with this property?
Editors’ comment: A quadratic function with that property was proved to exist in the special case where $G=H\rtimes C$ and $C$ is a cyclic subgroup such that $C_{H}(C)=1$, (E. Jabara, J. Algebra, 487 (2017), 161–172). A. Turull
@@
18.104. Construct an example of a pro-$p$ group $G$ and a proper abstract normal subgroup $K$ such that $G/K$ is perfect. J. S. Wilson
@@
If so, then ${\rm SL}_{3}(R)$ and the kernel of the map to ${\rm SL}_{3}(R/U)$ provide an answer to the previous question. J. S. Wilson
@@
The motivation is the theorem in (E. Breuillard, B. Green, M. Kassabov, A. Lubotzky, N. Nikolov, T. Tao) saying that there are $\epsilon>0$ and $k$ such that every finite simple group $G$ contains a generating set $S$ with at most $k$ elements such that the Cayley graph of $G$ with respect to $S$ is an $\epsilon$-expander. In (M. Ershov, A. Jaikin-Zapirain, M. Kassabov, Mem. Amer. Math. Soc., 1186, 2017) it is proved that there exists a group with Kazhdan’s property (T) that maps onto every simple group of Lie type of rank $\geqslant 2$. (A group $G$ is said to have Kazhdan’s property (T) if its trivial representation is an isolated point in the unitary dual space of $G$.) A. Jaikin-Zapirain
@@
Classify the basic primitive groups $G$ which do not have the road closure property. In particular, is it true that such a group either has a subgroup of index at most 3 or is almost simple? P. J. Cameron
@@
The conjecture is proved if $\varphi$ has prime order (E. Jabara, J. Algebra, 320, no. 10 (2008), 3671–3679). If $G$ is infinitely generated, then $G$ does not have to be almost solvable (K. Dekimpe, D. Gonçalves, Bull. London Math. Soc., 46, no. 4 (2014), 737–746), but if it is of finite upper rank, then it has to be almost solvable (E. Troitsky, J. Group Theory, 28, no. 1 (2025), 151–164). A. L. Fel’shtyn, E. V. Troitsky
@@
19.31. Let $\omega(G,S)$ be the exponential growth rate of a group $G$ with a finite generating set $S$ (see 14.7). Let $MCG(\Sigma_{g})$ be the mapping class group of the orientable surface $\Sigma_{g}$ of genus $g$. Is there a constant $C>1$ such that $\omega(MCG(\Sigma_{g}),S)\geqslant C$ for every $g>0$ and every finite generating set $S$ of $MCG(\Sigma_{g})$? K. Fujiwara
No, there is no such constant that works for all genera; but if the genus $g$ is fixed, then such a constant $C=C(g)>1$ does exist (J. Mangahas, Geom. Funct. Anal. 19, no. 5 (2010), 1468–1480).